If \\(z_1=3+2i\\) and \\(z_2=4+3i\\) and are complex numbers, find \\(z_1z_2\\)
1
Expert's answer
2020-06-15T15:18:39-0400

z_1z_2 = (3+2i)(4+3i) = 12+8i+9i+6i^2= 12+17i-6=6+17i.

Answer: z_1z_2=6+17i.